OpenStudy (anonymous):

please help 2- a-10/a+4 = a^2-2/a+4

OpenStudy (anonymous):

\[\frac{2 - a -10}{a+4} = \frac{a^2 -2}{a+4} \]

OpenStudy (anonymous):

Is that what it looks like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, well the first thing I would do is try to factor the numerators to see if we can cancel anything

OpenStudy (anonymous):

except for 2- then the a-10/a+4= a^2-2/a+4

OpenStudy (anonymous):

If there is something that's in parentheses let me know

OpenStudy (anonymous):

no, its not in parentheses

OpenStudy (anonymous):

so the fraction is correct as written, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, hold on

OpenStudy (anonymous):

1. Okay, so basically the only thing that simplifies is 2 - a - 10. It simplifies to -a - 8. 2. We will need to cross multiply: \[\frac{-(a+8)}{a+4}=\frac{a^2-2}{a+4}\]

OpenStudy (anonymous):

3. The result of cross multiplying gives us: \[-(a+8)(a+4)=(a^2-2)(a+4)\]

OpenStudy (anonymous):

Now we will have to multiply everything back out. We can't divide both sides by (a+4) because it will eliminate one of the solutions for the variable a

OpenStudy (anonymous):

\[-a^2-12a-32 = a^3+4a^2-2a-8\]

OpenStudy (anonymous):

Now we will have to move like terms on one side:

OpenStudy (anonymous):

Put the variables on one side and put the last coefficients on the other side

OpenStudy (anonymous):

\[a^3 +3a^2 +10a = -24\]

OpenStudy (anonymous):

this is still somewhat confusing to me, but I am trying to understand it more

OpenStudy (anonymous):

Wait a minute..let me post twiddla. You can post how the problem looks on there

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

Okay, so it all came out to: \[a^2 - a - 20\] (a-5)(a+4)

OpenStudy (anonymous):

So the solution is a = 5, -4

OpenStudy (anonymous):

so -4 is the other solution for a

OpenStudy (anonymous):

lost connection to to other

OpenStudy (anonymous):

That's okay